OpenStudy (anonymous):

Kathy is 6 years younger than Bill. Twelve years ago, Bill was twice as Kathy was then. How old are they know?

OpenStudy (anonymous):

k = b - 6 2(k - 12) = b-12 2k -24 = b-12 2k = b-12+24 2k = b + 12 //substitute 2(b - 6) = b +12 2b - 12 = b +12 2b = b +24 b = 24 /substitute in original k = b - 6 k = 24-6 k = 18 24 and 18 12 years ago 24-12 and 18-12 or 12 and 6

OpenStudy (anonymous):

Let k and b be Kathy's and Bill's current age respectively. Solve the following equations for k and b:\[\left\{k=b-6, \frac{b-12}{k-12}=2\right\}\]\[\{k=18,b=24\} \]